OpenStudy (anonymous):

Which of the following is the simplified form of the quantity x squared minus 9 over the quantity x minus 3 ? x – 3, with the restriction x ≠ 3 x + 3, with the restriction x ≠ 3 x – 3, with the restriction x ≠ –3 x + 3, with the restriction x ≠ –3

OpenStudy (luigi0210):

you got this! >.<

OpenStudy (anonymous):

I don't I'm stupid... lol

OpenStudy (luigi0210):

No you're not.

OpenStudy (anonymous):

Feels like it, is it possible for someone to have creeper status on this thing? Like IDK a stalker?

OpenStudy (luigi0210):

Hm, probably, why?

OpenStudy (anonymous):

I think I may have one.... any who. Do you know how to do this?

OpenStudy (anonymous):

Yup. I got a stalker/rapist.

OpenStudy (luigi0210):

Ha, naw :P Write out the problem.

OpenStudy (anonymous):

Okay,

OpenStudy (anonymous):

\[\frac{ x^2-9 }{ x-3 }\]

OpenStudy (luigi0210):

You can factor the top equation right?

OpenStudy (anonymous):

I think... i'm not suree..

OpenStudy (luigi0210):

Well try it :)

OpenStudy (anonymous):

I don't know if it be right,

OpenStudy (anonymous):

But the top part I would just say it was 8x

OpenStudy (luigi0210):

8x?

OpenStudy (anonymous):

Yeah, I'm not sure what to do.. like I said I'm stupid.

OpenStudy (luigi0210):

No, you are not stupid! >.<

OpenStudy (anonymous):

Would it be B?

OpenStudy (luigi0210):

o_o Lucky guess or did you work it out?

OpenStudy (anonymous):

I just did -9 divided by -3

OpenStudy (luigi0210):

\[\frac{ x^2-9 }{ x-3 }=\frac{ (x-3)(x+3) }{ x-3 }\]

OpenStudy (luigi0210):

Notice anything that can cancel?

OpenStudy (anonymous):

negative 3

OpenStudy (luigi0210):

yup, so that leaves x+3, with a restriction of\[x \neq 3\]

OpenStudy (anonymous):

Ah, well the rest of them aren't that easy...

OpenStudy (luigi0210):

x-3=0 +3 +3 --------------- x=3

OpenStudy (luigi0210):

Don't worry, I got you :) and anyone else on OS :P

OpenStudy (anonymous):

Thanks (:

OpenStudy (luigi0210):

You're welcome :D